OpenStudy (anonymous):

The Flemings secured a bank loan of $248,000 to help finance the purchase of a house. The bank charges interest at a rate of 10%/year on the unpaid balance, and interest computations are made at the end of each month. The Flemings have agreed to repay the loan in equal monthly installments over 25 years. What should be the size of each repayment if the loan is to be amortized at the end of the term? (Round answer to the nearest cent.)

OpenStudy (tkhunny):

I'd try this. \(248000 = Pmt\cdot\dfrac{1 - v^{25*12}}{i}\) Where i = 0.10/12 and v = 1/(1+i)

OpenStudy (anonymous):

It worked, thank you.